OpenStudy (anonymous):

the solution to the differential equation dy/dx=10xy with the initial condition y(0)=2 is?

OpenStudy (eyust707):

I believe you can just separate the variables

OpenStudy (eyust707):

then integrate both side

OpenStudy (eyust707):

youll end up with a constant from the integration

OpenStudy (anonymous):

would i take out the 10?

OpenStudy (eyust707):

no

OpenStudy (eyust707):

\[(1/y)9 dy = 10x dx\]

OpenStudy (anonymous):

oh okay. so the 10 goes with the x? and the answer choices all have ln or e in them

OpenStudy (eyust707):

\[\int\limits (1/y) dy = \int\limits 10xdx\]

OpenStudy (eyust707):

the 10 could go on either side but with the x is much easier

OpenStudy (anonymous):

okay so far i have ln y=5x^2

OpenStudy (anonymous):

That's correct :)

OpenStudy (anonymous):

Should add C, though!

OpenStudy (anonymous):

according to the answer choices i have, im done yet lol

OpenStudy (anonymous):

You need to find y!

OpenStudy (anonymous):

yeah i guess. then i got y=5x^2/ln, which should equal e^5x^2 right?

OpenStudy (anonymous):

Yep, still don't forget C

OpenStudy (anonymous):

then if i plugged in y(0)=2 i would get e^5x^2?

OpenStudy (anonymous):

the answer choices are A. ln((5x^2)+2) b.e^5x^2+2 C. e^5x^2+1 D.2 ln(5x^2) E. 2e^5x^2

OpenStudy (anonymous):

What did I keep reminding you ? The constant C!

OpenStudy (anonymous):

ohhh....i got you know. the constant c would be 2. oh wow i feel stupid now

OpenStudy (anonymous):

can you help me with another question?

OpenStudy (anonymous):

:) So you know the purpose of initial value now!

OpenStudy (anonymous):

Just open the new post!

OpenStudy (anonymous):

okay i did. and thank you for your assistance with the last problem